OpenStudy (anonymous):

A survey of 1,000 men and women asked, "Do you earn over $50,000 per year?" The table below shows the responses for males and females.

OpenStudy (anonymous):

Male Female Total Income over $50,000 485 385 870 Income below $50,000 65 65 130 Total 550 450 1,000

OpenStudy (anonymous):

Based on these data, are "being male" and "earning over $50,000" independent events? (5 points) Yes, P(being male | the person earns over $50,000) ≠ P(being male) Yes, P(being male | the person earns over $50,000) = P(being male) No, P(being male | the person earns over $50,000) ≠ P(being male) No, P(being male | the person earns over $50,000) = P(being male)

OpenStudy (anonymous):

Is it the last one? @ranga

OpenStudy (kropot72):

A and B are independent if P(A|B) = P(A). In this case A is 'being male', and B is 'the person earns over $50,000'. P(being male) = 550/1000 = 0.55 P(being male | the person earns over $50,000) = 485/870 = 0.557 So we find: No, P(being male | the person earns over $50,000) ≠ P(being male)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (kropot72):

You're welcome :)